write the first five terms of the sequence defined by the given recursive or explicit formula. tn=32-8n
Answer by Edwin McCravy(20086)

write the first five terms of the sequence defined by the given recursive or explicit formula.
tn=32-8n


This is an explicit formula because it does not have tn on
one side and tn+1 or tn-1 on the other, and
it does not give separately the value of the first term.

tn=32-8n

To find the first term, t1, we substitute 1 for n

tn=32-8n

t1=32-8(1)

t1=32-8 

t1=24

To find the second term, t2, we substitute 2 for n

tn=32-8n

t2=32-8(2)

t2=32-16 

t2=16

To find the third term, t3, we substitute 3 for n

tn=32-8n

t3=32-8(3)

t3=32-24 

t3=8

To find the fourth term, t4, we substitute 4 for n

tn=32-8n

t4=32-8(4)

t4=32-32 

t4=0

To find the fifth term, t5, we substitute 5 for n

tn=32-8n

t5=32-8(5)

t5=32-40 

t5=-8

So the 5 terms of the sequence are 24, 16, 8, 0, -8 

-------------------------------

You didn't ask this, but you'll need to know it:

The same sequence could have been given to you 
recursively as:

tn+1 = tn-8, t1 = 24

Notice that it has t's on both sides of the equation and
the first term t1 is given.

Then we would substitute 2 for n and 24 for t1 to find t2

tn+1 = tn-8
t1+1 = t1-8
t2 = t1-8
t2 = 24-8
t2 = 16

Then we would substitute 3 for n and 16 for t2 to find t3

tn+1 = tn-8
t2+1 = t2-8
t3 = t2-8
t3 = 16-8
t3 = 8

Then we would substitute 4 for n and 8 for t3 to find t4

tn+1 = tn-8
t3+1 = t3-8
t4 = t3-8
t4 = 8-8
t4 = 0

Then we would substitute 5 for n and 0 for t4 to find t5

tn+1 = tn-8
t4+1 = t4-8
t5 = t4-8
t5 = 0-8
t5 = -8

So the 5 terms of the sequence are 24, 16, 8, 0, -8 

Edwin